OpenStudy (anonymous):

the numerical value of 15 sec 20° cos 20° is

OpenStudy (campbell_st):

well sec = 1/cos so you have \[15 \times \frac{1}{\cos(20)} \times \cos(20).... or... 15 \times \frac{\cos(20)}{\cos(20)}\] I'll leave you to finish it